OpenStudy (anonymous):

Fraction 4 over 5n = Fraction 2 over 3. n = ___? Fraction 2 over 15 Fraction 5 over 6 1Fraction 1 over 5 1Fraction 7 over 15 @iGreen

OpenStudy (igreen):

\(\dfrac{4}{5n} = \dfrac{2}{3}\) Cross multiply. 4 * 3 = 5n * 2

OpenStudy (anonymous):

i still dont get it @iGreen